OpenStudy (anonymous):

Linear equations 3x-8y=7 9x-24y=35 is the answer No Solution?

OpenStudy (anonymous):

Okay. you have 3x-8y=7 and you have 9x-24y=35 you take 3x-8y=7 out of standard form by subtracting 3x from both sides. once you do that it becomes -8y=-3x+7. (Because they are not like terms you leave them separate, just on the same side of the equal sign.) Then you divide your -8 to both sides to get y by it's self. so it becomes slope intercept form. Leaving you with y=3/8x-7/8 Then do the same with the other one and you get your answer. Medal? If helped?

OpenStudy (anonymous):

So the answer is y= 5/4 and x=17/3 ?